OpenStudy (anonymous):

Leslie set up a fruit stand in order to sell some of her crop of oranges. the first person who came by bought 1/3 of her oranges. the second person bought four. the third person then bought 1/4 of the remaining oranges. leslie took the last 15 oranges home and made orange juice. how many oranges did leslie have at the beginning

OpenStudy (anonymous):

Let x=number of oranges that Leslie had at the beginning After the first person came and bought a third, she had x-(1/3)x=(2/3)x left After the second person bought 4, she had (2/3)x-4 left After the third person bought a fourth of the remaining, she had ((2/3)x-4))-1/4((2/3)x-4) left and we are told this equals 15, sooooo ((2/3)x-4)-1/4((2/3)x-4)=15 lets first multiply each term by 4 4(2/3)x-4)-(2/3)x+4=60 (8/3)x-16-(2/3)x+4=60 (6/3)x-12=60 2x=72 x=36----amount she had at the beginning Ck 36-(1/3)*36=24 24-4=20 20-(1/4)*20=15 15=15